# Cascaded Large-Scale TSP Solving with Unified Neural Guidance: Bridging Local and Population-based Search

### v1

@@ 4.2 Results and Analysis
Table 2: Optimality gap (%) on the Ex-Large set (10 runs per instance). “-” in Best / Avg columns indicates no solutions produced in 10 runs / solutions produced only in some runs. Bold indicates best performance. “Total” row summarizes average gaps across Ex-Large set.
@@ 4.3 Ablation Study
Phase Transition Policy. The effectiveness of the learning-based transition policy within UNG was evaluated by comparing it against fixed transition times. During policy training, $t_{trans}$ varied from 50s to 650s. In this study, fixed $t_{trans}$ values of 50s, 200s, 350s, 500s, and 650s were tested while maintaining all other UNiCS settings. Figure 5(a) and Figure 5(b) illustrate the resultant convergence curves on Large and Ex-large sets. It can be observed that when $t_{trans}$ is fixed at a small value (e.g., 50s), UNiCS’s performance on the Large set closely matches that of UNiCS using the transition policy. However, as $t_{trans}$ increases, the performance with fixed transition times noticeably deteriorates. Conversely, Figure 5(b) demonstrates the opposite pattern: with large fixed $t_{trans}$ values (e.g., 650s), performance on the Ex-large set matches that of the transition policy, but deteriorates significantly as $t_{trans}$ decreases. These results indicate that fixed $t_{trans}$ fails to adapt to instances of varying sizes, while the transition policy maintains strong performance across different instances. A scatter plot of the collected data for training the transition policy is provided in Appendix B, from which an approximately linear relationship between problem size and optimal transition time can be observed.
Neural Guidance for $AB$-cycle Selection. Previous results on both Large and Ex-large sets demonstrates that UNiCS-P typically achieves better performance than the original EAX method. To further validate the effectiveness of UNG guidance for $AB$-cycle selection within UNiCS’s cascaded framework, an ablation study was conducted by replacing PBS phase in UNiCS with the original EAX method. The resultant method is referred to as UNiCS*. Figure 5(c) shows convergence curves on Ex-large set (complete results provided in Appendix A.3). The results show that UNiCS* converges significantly slower than UNiCS, confirming the effectiveness of neural guidance for $AB$-cycle selection within the cascaded framework. Meanwhile, the fact that UNiCS* outperforms both EAX and UNiCS-P demonstrates the inherent benefits of combining the strengths of LS and PBS.
@@ 5 Conclusion
This work introduces UNiCS, which employs a UNG module to direct both LS and PBS, effectively combining their complementary strengths for solving large-scale TSP instances. With UNG guiding solution generation and determining phase transition timing, UNiCS demonstrates superior solution quality compared to state-of-the-art heuristic and hybrid methods across various runtime budgets. The strong performance on challenging benchmarks, despite being trained only on simple distributions, demonstrates UNiCS’s generalization capabilities. One limitation of the present work is that, while problem size is a crucial factor in identifying appropriate transition time, it may not fully determine the optimal transition time for instances beyond those tested in this work. Future research will explore incorporating additional instance characteristics to develop more advanced transition policies.
@@ A.1.1 TSP Benchmarks
Figure 8: Cumulative optimality gaps on Ex-Large set with instances sorted by ascending problem size..
@@ A.1.2 Compared Methods
For LKH and VSR-LKH, in addition to the original implementation using $\alpha$-measure method, a variant using POPMUSIC Helsgaun [2018] was also evaluated.
@@ A.2.3 Results on each TSP instance:
Tables 3-8 (at the end of the Appendix) present the performance of all solvers on each instance at different solving times, respectively. It can be observed that the UNiCS not only achieves the best performance in terms of the overall average gap but also demonstrates superior performance on larger instances overall.
@@ B.1 Training Detail
Figure 13: Convergence curves for different values of $t_{trans}$ on pbh30400.
@@ Appendix C The EAX Method
  • In Stage 2, the Block2 Strategy is employed, which focuses on constructing an E-set with relatively few and large segments of the parent solutions ($p_{A}$ and $p_{B}$). This strategy aims to reduce the number of subtours in the intermediate solutions, thereby increasing the likelihood of generating high-quality offspring. The Block2 Strategy involves a tabu search process to efficiently explore combinations of $AB$-cycles, ensuring that the resulting $E$-set is effective in producing improved solutions. The detailed mechanism of the Block2 Strategy, including the use of central $AB$-cycles and the calculation of #$\#C$ values to minimize subtours, is crucial in the later stages of the search, where more significant structural changes are necessary to escape local optima.

### v2

@@ 4.2 Results and Analysis
Table 2: Relative gap (%) to the BKS on the Ex-Large set (10 runs per instance). “-” in Best / Avg columns indicates no solutions produced in 10 runs / solutions produced only in some runs. Bold indicates best performance. “Total” row summarizes average gaps across Ex-Large set.
@@ 4.3 Ablation Study
Phase Transition Policy. The effectiveness of the learning-based transition policy within UNG was evaluated by comparing it against fixed transition times. During policy training, $t_{trans}$ varied from 50s to 650s. In this study, fixed $t_{trans}$ values of 50s, 200s, 350s, 500s, and 650s were tested while maintaining all other UNiCS settings. Figure 5(a) and Figure 5(b) illustrate the resultant convergence curves on Large and Ex-large sets. It can be observed that when $t_{trans}$ is fixed at a small value (e.g., 50s), UNiCS’s performance on the Large set closely matches that of UNiCS using the transition policy. However, as $t_{trans}$ increases, the performance with fixed transition times noticeably deteriorates. Conversely, Figure 5(b) demonstrates the opposite pattern: with large fixed $t_{trans}$ values (e.g., 650s), performance on the Ex-large set matches that of the transition policy, but deteriorates significantly as $t_{trans}$ decreases. These results indicate that fixed $t_{trans}$ fails to adapt to instances of varying sizes, while the transition policy maintains strong performance across different instances. A scatter plot of the collected data for training the transition policy is provided in Section B in [20], from which an approximately linear relationship between problem size and optimal transition time can be observed.
Neural Guidance for $AB$-cycle Selection. Previous results on both Large and Ex-large sets demonstrates that UNiCS-P typically achieves better performance than the original EAX method. To further validate the effectiveness of UNG guidance for $AB$-cycle selection within UNiCS’s cascaded framework, an ablation study was conducted by replacing PBS phase in UNiCS with the original EAX method. The resultant method is referred to as UNiCS*. Figure 5(c) shows convergence curves on Ex-large set. The results show that UNiCS* converges significantly slower than UNiCS, confirming the effectiveness of neural guidance for $AB$-cycle selection within the cascaded framework. Meanwhile, the fact that UNiCS* outperforms both EAX and UNiCS-P demonstrates the inherent benefits of combining the strengths of LS and PBS.
@@ 5 Conclusion
This work introduces UNiCS, a solver that employs a Unified Neural Guidance (UNG) module to effectively combine the complementary strengths of Local Search (LS) and Population-Based Search (PBS). By guiding solution generation and phase transition timing, UNiCS consistently outperforms state-of-the-art methods on large-scale TSP instances across various runtime budgets. Strong performance on challenging benchmarks demonstrates its generalization capabilities, while ablation studies further validate the effectiveness of UNG in guiding AB-cycle selection and phase transition.
While the transition policy in UNiCS is simple and effective, its adaptability can be further improved. A key limitation is its reliance on problem size, which may not be optimal for all instances. Future research will focus on developing a more adaptive policy by incorporating additional instance characteristics (e.g., node distribution) and hardware configurations, thereby enhancing UNiCS’s instance and hardware-agnostic generalization. Additionally, with the rise of Large Language Models (LLMs), recent work has demonstrated their potential for guiding optimization solvers[19], which also represents a promising direction for our future research.
{ack}
This work was supported by National Key Research and Development Program of China under Grant 2022YFA1004102, in part by Zhongguancun Academy Project No.20240303, in part by National Natural Science Foundation of China under Grant 72401105, and in part by Hubei Provincial Natural Science Foundation of China under Grant 2024AFB338.
@@ A.1.1 TSP Benchmarks
Figure 8: Cumulative optimality gaps on Ex-Large set with instances sorted by ascending problem size.
@@ A.1.2 Compared Methods
For LKH and VSR-LKH, in addition to the original implementation using $\alpha$-measure method, a variant using POPMUSIC [12] was also evaluated.
@@ A.2.3 Results on each TSP instance:
Tables 6-11 (at the end of the Appendix) present the performance of all solvers on each instance at different solving times, respectively. It can be observed that the UNiCS not only achieves the best performance in terms of the overall average gap but also demonstrates superior performance on larger instances overall.
@@ B.1 Training Detail
Figure 13: Convergence curves for different values of $t_{\text{trans}}$ on pbh30400.
@@ Appendix C The EAX Method
  • In Stage 2, the Block2 Strategy is employed, which focuses on constructing an E-set with relatively few and large segments of the parent solutions ($p_{A}$ and $p_{B}$). This strategy aims to reduce the number of subtours in the intermediate solutions, thereby increasing the likelihood of generating high-quality offspring. The Block2 Strategy involves a tabu search process to efficiently explore combinations of $AB$-cycles, ensuring that the resulting $E$-set is effective in producing improved solutions. The detailed mechanism of the Block2 Strategy, including the use of central $AB$-cycles and the calculation of #C\#C values to minimize subtours, is crucial in the later stages of the search, where more significant structural changes are necessary to escape local optima.
@@ D.2 The Original LKH Method
Appendix E Comparison with Deep Learning-based Methods
E.1 TSP Benchmarks
Due to the instance size-dependent GPU memory consumption and inference time of LEHD, 33 instances smaller than 10,000 nodes were selected from the TSPLib, National, and VLSI benchmark sets. These instances were divided into two subsets: the Small set (15 instances, 3,500–5,000 nodes) and the Medium set (18 instances, 5,000–10,000 nodes). Due to the non-configurable and relatively short runtime nature of LEHD, only the Local Search (LS) phase was tested for comparison, denoted as UNiCS-LS.
E.2 Testing Method
The same hardware configuration as in the main experiments was employed: two AMD EPYC 7713 CPUs and a NVIDIA A800 80GB. For each instance, all methods were executed with 10 independent runs, following the protocol of the main experiments. To ensure fair comparisons, runtime parity between LEHD and UNiCS-LS was enforced. Because LEHD does not support explicit runtime configuration, the following procedure was adopted: LEHD was first executed to record its actual runtime per instance, and UNiCS-LS was then constrained to the identical runtime for that instance.
Table 3: Relative gap (%) to the BKS on the Large set (10 runs per instance). “-” in Best / Avg columns indicates no solutions produced in 10 runs / solutions produced only in some runs. Bold indicates best performance. “Total” row summarizes average gaps across Large set.
Benchmark UNiCS-LS LEHD
Best Avg Best Avg
Small,(3500-5000] 0.1706 0.2848 22.4353 22.4353
Medium,(5000-10000] 0.0246 0.0689 62.6516 62.6516
E.3 Results Analysis
Table 3 summarizes the results on the Small and Medium datasets. LEHD, being a deterministic method, yields identical Best and Avg results. As shown, LEHD exhibits performance gaps of orders of magnitude compared to UNiCS-LS on both subsets, with even larger discrepancies on Medium. This indicates that neural network-based methods like LEHD remain inadequate for solving large-scale TSP instances derived from real-world scenarios. Notably, the Medium instances are still smaller than the smallest instances in our main experiments (Large set), justifying our omission of LEHD from the main text. For more detailed results, refer to Table 4 and Table 5.
While other deep learning-based methods (as reported in their papers) show no orders-of-magnitude differences from LEHD on TSPLib, our method achieves significantly better results than LEHD on Medium. These findings suggest that neural methods and our approach differ fundamentally in focus and performance: neural methods excel in small-scale instances (e.g., <1,000 nodes) with shorter runtimes but struggle with scalability. Thus, direct comparisons between our method and neural approaches are inappropriate, as their strengths lie in distinct problem regimes.
Table 4: Relative gap (%) to the BKS on the Small set (10 runs per instance).
Instance Optimum Run Time(s) UNiCS-LS LEHD
Best(Gap%) Average(Gap%) Best(Gap%) Average(Gap%)
fjs3649 9272.0 39.69 9272.0(0.0000) 9273.4(0.0151) 12862.13281(38.7200) 12862.13281(38.7200)
fjr3672 9601.0 40.31 9601.0(0.0000) 9601.8(0.0083) 12959.99414(34.9900) 12959.99414(34.9900)
dlb3694 10959.0 41.27 10959.0(0.0000) 10960.4(0.0128) 12596.28711(14.9400) 12596.28711(14.9400)
ltb3729 11821.0 41.75 11822.0(0.0085) 11824.6(0.0305) 13809.67871(16.8200) 13809.67871(16.8200)
fl3795 28772.0 43.35 29465.0(2.4086) 29872.4(3.8246) 34838.125(21.0800) 34838.125(21.0800)
xqe3891 11995.0 46.88 12000.0(0.0417) 12005.8(0.0900) 13756.92188(14.6900) 13756.92188(14.6900)
xua3937 11239.0 48.29 11240.0(0.0089) 11241.2(0.0196) 12669.81055(12.7300) 12669.81055(12.7300)
dkc3938 12503.0 47.02 12508.0(0.0400) 12510.8(0.0624) 15167.35547(21.3100) 15167.35547(21.3100)
dkf3954 12538.0 48.39 12538.0(0.0000) 12545.0(0.0558) 14042.00879(12.0000) 14042.00879(12.0000)
bgd4396 13009.0 63.5 13010.0(0.0077) 13015.8(0.0523) 15522.55078(19.3200) 15522.55078(19.3200)
frv4410 10711.0 62.91 10713.0(0.0187) 10715.8(0.0448) 13143.87891(22.7100) 13143.87891(22.7100)
fnl4461 182566.0 64.86 182566.0(0.0000) 182569.0(0.0016) 206228.8594(12.9600) 206228.8594(12.9600)
bgf4475 13221.0 67.69 13224.0(0.0227) 13226.8(0.0439) 16032.89941(21.2700) 16032.89941(21.2700)
ca4663 1290319.0 72.32 1290355.0(0.0028) 1290388.6(0.0054) 1705001.25(32.1400) 1705001.25(32.1400)
xqd4966 15316.0 82.91 15316.0(0.0000) 15316.8(0.0052) 21572.55859(40.8500) 21572.55859(40.8500)
Avg Avg Avg 0.1706 0.2848 22.4353 22.4353
Table 5: Relative gap (%) to the BKS on the Small set (10 runs per instance).
Instance Optimum Run Time(s) UNiCS-LS LEHD
Best(Gap%) Average(Gap%) Best(Gap%) Average(Gap%)
fqm5087 13029.0 85.25 13035.0(0.0461) 13032.0(0.0230) 18167.37891(39.4400) 18167.37891(39.4400)
fea5557 15445.0 108.42 15449.6(0.0298) 15445.0(0.0000) 18359.70313(18.8700) 18359.70313(18.8700)
rl5915 565530.0 127.29 565651.8(0.0215) 565615.0(0.0150) 672688.875(18.9500) 672688.875(18.9500)
rl5934 556045.0 128.22 556595.8(0.0991) 556152.0(0.0192) 648011.4375(16.5400) 648011.4375(16.5400)
tz6117 394718.0 136.84 394919.2(0.0510) 394751.0(0.0084) 599770.875(51.9500) 599770.875(51.9500)
xsc6880 21535.0 187.09 21544.8(0.0455) 21543.0(0.0371) 25774.41016(19.6900) 25774.41016(19.6900)
eg7146 172386.0 204.8 173139.8(0.4373) 172739.0(0.2048) 281994.0313(63.5800) 281994.0313(63.5800)
bnd7168 21834.0 210.08 21838.0(0.0183) 21834.0(0.0000) 26495.96484(21.3500) 26495.96484(21.3500)
lap7454 19535.0 234.16 19541.4(0.0328) 19537.0(0.0102) 23547.13672(20.5400) 23547.13672(20.5400)
ym7663 238314.0 248.69 238321.4(0.0031) 238317.0(0.0013) 448895.8438(88.3600) 448895.8438(88.3600)
pm8079 114855.0 286.15 115128.8(0.2384) 114931.0(0.0662) 329506.5938(186.8900) 329506.5938(186.8900)
ida8197 22338.0 303.45 22344.8(0.0304) 22338.0(0.0000) 27432.51172(22.8100) 27432.51172(22.8100)
ei8246 206171.0 302.44 206177.8(0.0033) 206173.0(0.0010) 615127.375(198.3600) 615127.375(198.3600)
ar9152 837479.0 406.42 838056.2(0.0689) 837768.0(0.0345) 1281666.75(53.0400) 1281666.75(53.0400)
dga9698 27724.0 482.04 27728.8(0.0173) 27727.0(0.0108) 34786.00391(25.4700) 34786.00391(25.4700)
ja9847 491924.0 494.19 492377.6(0.0922) 491970.0(0.0094) 1255864.75(155.3000) 1255864.75(155.3000)
gr9882 300899.0 499.78 300903.0(0.0013) 300899.0(0.0000) 549525.125(82.6300) 549525.125(82.6300)
kz9976 1061881.0 520.35 1061922.6(0.0039) 1061891.0(0.0009) 1528685.0(43.9600) 1528685.0(43.9600)
Avg Avg Avg 0.0246 0.0689 62.6516 62.6516